OpenStudy (mellamokatie):

Explain how to write a function rule from the table below. Then write a function rule. x|2|4| 6 y|1|0|-1

OpenStudy (mellamokatie):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

notice how x increases by 2 each time what is y decreasing by each time?

OpenStudy (mellamokatie):

because they're polar opposite's?

jimthompson5910 (jim_thompson5910):

I don't know what you mean here

OpenStudy (mellamokatie):

That they are opposite's? I dont know

jimthompson5910 (jim_thompson5910):

going from y = 1 to y = 0 is a decrease of 1, agreed?

OpenStudy (mellamokatie):

yeah

jimthompson5910 (jim_thompson5910):

going from y = 0 to y = -1 is also a decrease of 1, right?

OpenStudy (mellamokatie):

yes

jimthompson5910 (jim_thompson5910):

so because y goes down the same amount each time, this suggests we have a linear equation

jimthompson5910 (jim_thompson5910):

x goes up by 2, y goes down by 1 slope = rise/run = -1/2 = -0.5 agreed? or no?

OpenStudy (mellamokatie):

Oh okay this makes sense

jimthompson5910 (jim_thompson5910):

now pick any point from the table, say (x,y) = (4,0) use m = -0.5 x = 4 y = 0 to find the equation of the line y = mx+b y = -0.5*x+b 0 = -0.5*4+b solve for b